144. Touchdown. Suppose at the kickoff of a football at the left side of the goal line and runs for a touchdown diagonally across the field. How many yards would he run? (A football field is 100 yard long and 160 feet wide.)
Answer by checkley71(8403)

USING THE PATHEGOREAN THEROM A^2+B^2=C^2 WE GET
300^2+160^2=C^2
90000+25600=C^2
C^2=115600
C=SQRT115600
C=340 FEET OR 340/3=113 YARDS & 1 FOOT (113.3333YDS.)IS THE DIAGONAL ACROSS THE FOOTBALL FIELD.